# Multi-user QKD using quotient graph states derived from continuous-variable dual-rail cluster states

### v1

@@ Abstract
Multipartite entangled states are fundamental resources for multi-user quantum cryptographic tasks. Despite significant advancements in generating large-scale continuous-variable (CV) cluster states, particularly the dual-rail cluster state because of its utility in measurement-based quantum computation, its application in quantum cryptography has remained largely unexplored. In this paper, we introduce a novel protocol for generating three user conference keys using a CV dual-rail cluster state. We develop the concept of a quotient graph state by applying a node coloring scheme to the infinite dual-rail graph, resulting in a six-mode pure graph state suitable for cryptographic applications. Our results demonstrate that the proposed protocol achieves performance close to that of GHZ-based protocols for quantum conference key agreement (QCKA), with GHZ states performing slightly better. However, a key advantage of our protocol lies in its ability to generate bipartite keys post-QCKA, a feature not achievable with GHZ states. Additionally, compared to a downstream access network using two-mode squeezed vacuum states, our protocol achieves superior performance in generating bipartite keys. Furthermore, we extend our analysis to the finite-size regime and consider the impact of using impure squeezed states for generating the multipartite entangled states, reflecting experimental imperfections. Our findings indicate that even with finite resources and non-ideal state preparation, the proposed protocol maintains its advantages. We also introduce a more accurate method to estimate the capacity of a protocol to generate bipartite keys in a quantum network.
keywords:
CV-QKD, dual-rail cluster state, conference key agreement, GHZ state, quotient graph state
1 Introduction
@@
In the realm of CV quantum systems, the development and application of multipartite entangled states have predominantly focused on measurement-based quantum computation, particularly using cluster states [11, 12, 13]. Despite significant technological advances in the generation of large-scale CV cluster states, especially the dual-rail cluster state, these resources have not been thoroughly explored for quantum cryptographic applications. This gap is notable given the experimental maturity achieved in creating such states through techniques like time [14, 15, 16, 17] and frequency domain multiplexing [18, 19, 20, 21].
@@
Additionally, we consider the impact of using impure squeezed states for generating the multipartite entangled states. In practical scenarios, experimental imperfections lead to squeezed states that are not perfectly pure [24], affecting the overall performance of quantum protocols. By incorporating impure squeezed states into our analysis, we provide a more realistic assessment of our protocol’s robustness against imperfections in state preparation. We find that noisy squeezed states, when trusted and not attributed to eavesdropping, do not adversely affect the security of the protocol and may actually be advantageous. Our findings indicate that the proposed protocol remains effective even with finite resources and non-ideal state preparation.
@@
This paper is structured as follows: In Section 2, we delve into the complexities of classical information processing for multiple users. Section 3 presents the security analysis of a general multipartite state distributed among trusted users and provides their key rate expressions. In Section 4 , we discuss the generation of various tripartite states from squeezed states, including the impact of impurity. In Section 5, we introduce techniques for utilizing dual-rail cluster states in various cryptographic protocols. In section 6 we introduce and analysis various multi-user cryptographic protocols. In section 7 we summarize the major results of our analysis. Finally, in Section 8 , we review the outcomes of our research and discuss potential avenues for future work.
2 Multivariate Gaussian Correlations and Multi-User QKD
This section simplifies the discussion of multivariate Gaussian correlations by focusing on three users, A, B, and C, although the principles discussed can be extended to larger networks.
@@ 2 Multivariate Gaussian Correlations and Multi-User QKD
  • Conference Key Generation: User A shares the syndrome, enabling users B and C to establish a conference key.
@@
Elaborating on the second and third scenarios, consider three users, specifically $A,B$ and $C$. $A$ aims to generate a key with $B$ and $C$, which can be accomplished through two separate techniques. The first technique requires $A$ to convey the essential information for error correction to $B$ and $C$. Conversely, $B$ and $C$ could forward the necessary information for error correction to $A$. In the initial scenario, the keys generated will be perfectly correlated as both $B$ and $C$ use the data possessed by $A$ to generate a key. However, in the second scenario, the situation gets more complicated if $B$ and $C$ are correlated, resulting in a potential partial correlation of the keys. In the event of correlations, $B$ and $C$ cannot be treated separately, as the mutual information $I(A:B,C)\neq I(A:B)+I(A:C)$, but rather $I(A:B,C)=I(A:B)+I(A:C|B)=I(A:C)+I(A:B|C)$. The mutual information is defined as $I(i:j)=H(i)-H(i|j)=H(j)-H(j|i)$, and the conditional mutual information $I(i:j|k)=H(i|k)-H(i|j,k)$. $H(.)$ represents the Shannon entropy. When dealing with multiple correlated users, it becomes essential to ascertain the optimal syndrome length, that needs to be transferred for successful error correction. In the given example, $B$ and $C$ must jointly transmit an amount of information quantified as $H(BC|A)$ to $A$. The restriction lies in the collective syndrome delivered by $B$ and $C$. The individual syndromes of $B$, say $S_{b}$, and $C$, say $S_{c}$, can exceed in length, provided that $H(S_{B},S_{C})=H(BC|A)$.
@@
3 Security analysis
3.1 Holevo Bound
@@
In the context of CV-QKD, the Holevo bound represents the upper limit on the information accessible to an eavesdropper about the measurement outcomes of the parties whose data are used to generate a secure key. To determine this bound, we assume that the eavesdropper, $E$, is entangled with all the other parties involved in the protocol, denoted as $A$, $B$, and $C$ (we consider three parties for simplicity, but the underlying principles can be extended to more parties and modes), This entanglement leads us to posit that the entire system comprising the eavesdropper and the legitimate parties is in a pure state. Consequently, the von Neumann entropy of the combined system, $S(ABCE)$, is zero. Using the subadditivity property of von Neumann entropy, we obtain the relation $S(E)\leq S(ABC)$ [25]. Similarly, we have $S(E|A)\leq S(BC|A)$. Eve’s knowledge of the measurement outcome of $A$ is upper bounded by
$\chi_{A}=S(ABC)-S(BC|A).$
After $A$ generates a key with other correlated parties, $B$ and $C$ share a bipartite state $BC|A$, which could be used to generate a key among them. The Holevo bounds in this case are
$\chi_{B|A}=S(BC|A)-S(C|BA)$
and
$\chi_{C|A}=S(BC|A)-S(B|CA).$
@@
$\Sigma_{ij}=\frac{1}{2}\left\langle\hat{x}_{i}\hat{x}_{j}+\hat{x}_{j}\hat{x}_{i}\right\rangle-\left\langle\hat{x}_{i}\right\rangle\left\langle\hat{x}_{j}\right\rangle,$
[…]
@@
$S=\sum_{i}\left[\left(\nu_{i}+\frac{1}{2}\right)\log\left(\nu_{i}+\frac{1}{2}\right)-\left(\nu_{i}-\frac{1}{2}\right)\log\left(\nu_{i}-\frac{1}{2}\right)\right].$
@@
where $\Sigma_{p}$ denotes the covariance matrix of the unmeasured modes of the multipartite state, $\Sigma_{q}$ represents the covariance matrix of the measured modes, $\Sigma_{pq}$ is the covariance matrix between the measured and unmeasured modes, and $X$ is defined as $Diag(1,0)$ for measurements in the x-quadrature and $MP$ in Moore-Penrose inverse.
@@
Numerous measurement strategies were explored. We have determined that homodyne measurement of the x-quadrature is the best option for all the multipartite states and protocols presented in this paper. Therefore, any further mention of measurement should be interpreted as homodyne measurement of the x-quadrature.
3.2 Key Rates
Having addressed the complexities of multi-user QKD in the contexts of information reconciliation and the estimation of the Holevo bound, we are now positioned to establish expressions for various key rates. For a three-user system, three distinct types of keys can be generated:
@@ 3.2 Key Rates
$\beta$ is the reconciliation efficiency and the superscript signifies the reference data of the user $i$ used to generate the key.
@@
Once the data possessed by user $i$ has been used to generate a conference key, the conditional data possessed by the remaining two users can still be utilized to generate a key between them. The key rate for these scenarios are given by:
@@
A user can generate independent bipartite keys with the users with whom it is correlated, provided the mutual information of this user (say $i$) with the rest of the users is higher than the mutual information among the other users (ideally zero). The key rates are:
@@
and $K^{k}(i:k|j)$. In the scenario where the mutual information between the other users is zero, the equation simplifies to $K^{k}(i:k)$. Notice that even when $j$ and $k$ are identical, the order of key generation influences the key rates. It is more useful to consider the sum of independent keys that $i$ can generate with other users. The actual key rates with each user depend on the information disclosed for error correction, specifically the syndrome length. The sum of independent keys is given by:
$K^{j,k}(i:jk)=H(j,k)-\zeta H(j,k|i)-\chi_{jk}=K^{j}(i:j)+K^{k}(i:k|j)$ (6)
As discussed in Sec. 2, $H(j,k|i)$ represents the optimal amount of information that $j$ and $k$ need to communicate to $i$ to facilitate error correction and $\zeta\geq 1$ accounts for inefficient information reconciliation. For the symmetric case, the independent key rates can be computed as follows:
$K(i:j)=K(i:k)=\frac{1}{2}\left[H(j,k)-\zeta H(\mathbb{S}_{j},\mathbb{S}_{k})-S(ijk)+S(i\mid jk)\right]$ (7)
@@
3.3 Finite Size Analysis

### v2

# Multiuser QKD using quotient graph states derived from continuous-variable dual-rail cluster states

@@ Abstract
Multipartite entangled states are essential for multi-user quantum cryptography. While large-scale continuous-variable (CV) cluster states, particularly the dual-rail cluster state, are well-studied in measurement-based quantum computation, their cryptographic potential remains underexplored. Here, we propose a three-user conference key protocol using a CV dual-rail cluster state. By applying a node-coloring scheme to the infinite dual-rail graph, we create a six-mode pure graph state ideal for cryptographic tasks. Our results demonstrate near-GHZ performance for quantum conference key agreement (QCKA). Crucially, our protocol uniquely enables bipartite keys post-QCKA, which GHZ states cannot provide. It also surpasses two-mode squeezed vacuum states in generating bipartite keys within downstream access networks. Considering finite-size effects and impure squeezed states, our scheme remains robust despite experimental imperfections. We also introduce an enhanced method to more accurately estimate bipartite key generation capacity in quantum networks, paving the way for practical multi-user quantum cryptography.
I Introduction
@@
In the realm of CV quantum systems, the development and application of multipartite entangled states have predominantly focused on measurement-based quantum computation, particularly using cluster states [11, 12, 13, 14, 15]. Despite significant technological advances in the generation of large-scale CV cluster states, especially the dual-rail cluster state, these resources have not been thoroughly explored for quantum cryptographic applications. This gap is notable given the experimental maturity achieved in creating such states through techniques like time [16, 17, 18, 19] and frequency domain multiplexing [20, 21, 22, 23].
@@
The underlying concept of a ”quotient graph” is a well-established tool in graph theory [26] used to simplify complex graphs by partitioning vertices. This technique has found applications in quantum mechanics for studying symmetries in quantum graphs [27, 28, 29]. Our work extends this principle to a new domain: we apply the quotienting procedure not merely for analysis, but as a constructive method to engineer a target multipartite entangled state from a larger, periodic resource state. This approach retains the essential entanglement and connectivity properties required for cryptographic protocols.
@@
Additionally, we consider the impact of using impure squeezed states for generating the multipartite entangled states. In practical scenarios, experimental imperfections lead to squeezed states that are not perfectly pure [30], affecting the overall performance of quantum protocols. By incorporating impure squeezed states into our analysis, we provide a more realistic assessment of our protocol’s robustness against imperfections in state preparation.
@@
This paper is structured as follows: In Section II, we delve into the complexities of classical information processing for multiple users. Section III presents the security analysis of a general multipartite state distributed among trusted users and provides their key rate expressions. In Section IV , we discuss the generation of various tripartite states from squeezed states, including the impact of impurity. In Section V, we introduce techniques for utilizing dual-rail cluster states in various cryptographic protocols. In Section VI we introduce and analyze various multi-user cryptographic protocols. In Section VII we summarize the major results of our analysis. Finally, in Section VIII , we review the outcomes of our research and discuss potential avenues for future work.
II Multivariate Gaussian Correlations and Multi-User QKD
This section simplifies the discussion of multivariate Gaussian correlations by focusing on three users, $A$, $B$, and $C$, although the principles discussed can be extended to larger networks. To analyze the inter-user relationships and their implications for QKD, it is useful to classify the types of correlations that may occur. In the context of error correction and key generation, the structure of these correlations profoundly influences the strategies we must employ.
In what follows, we delineate three primary correlation scenarios that capture the essence of these interactions:
@@ II Multivariate Gaussian Correlations and Multi-User QKD
  • Conference Key Agreement (CKA): User A shares the syndrome, enabling users B and C to establish a conference key.
@@
Elaborating on the second and third scenarios, consider three users, specifically $A,B$ and $C$. AA aims to generate a key with $B$ and $C$, which can be accomplished through two separate techniques. The first technique requires $A$ to convey the essential information for error correction to $B$ and $C$. Conversely, $B$ and $C$ could forward the necessary information for error correction to $A$. In the initial scenario, the keys generated will be perfectly correlated as both $B$ and $C$ use the data possessed by $A$ to generate a key. However, in the second scenario, the situation gets more complicated if $B$ and $C$ are correlated, resulting in a potential partial correlation of the keys. In the event of correlations, $B$ and $C$ cannot be treated separately, as the mutual information $I(A:B,C)\neq I(A:B)+I(A:C)$, but rather $I(A:B,C)=I(A:B)+I(A:C|B)=I(A:C)+I(A:B|C)$. The mutual information is defined as I(i:j)=H(i)H(i|j)=H(j)H(j|i)I(i:j)=H(i)-H(i|j)=H(j)-H(j|i), and the conditional mutual information I(i:j|k)=H(i|k)H(i|j,k)I(i:j|k)=H(i|k)-H(i|j,k). $H(.)$ represents the Shannon entropy. When dealing with multiple correlated users, it becomes essential to ascertain the optimal syndrome length, that needs to be transferred for successful error correction. In the given example, $B$ and $C$ must jointly transmit an amount of information quantified as $H(BC|A)$ to $A$. The restriction lies in the collective syndrome delivered by $B$ and $C$. The individual syndromes of $B$, say $S_{B}$, and $C$, say $S_{C}$, can exceed in length, provided that $H(S_{B},S_{C})=H(BC|A)$.
@@
Fig. 1 illustrates the information reconciliation process for CKA with $A$ as the reference. In this scenario, after reconciliation, $B$ and $C$ can also extract a bipartite key (see parts (a) and (b) of Fig. 1). Fig. 2 depicts the information reconciliation process for independent bipartite key generation, where user $A$ establishes separate keys with $B$ and $C$.
Figure 1: (a) Information reconciliation for Conference Key Agreement (CKA) with user $A$ as the reference; (b) Bipartite information reconciliation post-CKA.
Figure 2: Information reconciliation for independent bipartite key generation between user $A$ and users $B$ and $C$. User AA establishes separate keys with $B$ and with $C$, each sending a syndrome $S_{B}$ or $S_{C}$ to $A$. The quantities $H(S_{B})$ and $H(S_{C})$ denote the entropies of the communicated syndromes, and their optimal joint entropy satisfies $H(S_{B},S_{C})=H(BC\mid A)$.
III Security analysis
III.1 Holevo bound
@@
In the context of CV-QKD, the Holevo bound represents the upper limit on the information accessible to an eavesdropper about the measurement outcomes of the parties whose data are used to generate a secure key. To determine this bound, we assume that the eavesdropper, $E$, purifies the state of the legitimate parties, denoted as $A$, $B$, and $C$ (we consider three parties for simplicity, but the underlying principles can be extended to more parties and modes) [31, 32]. This entanglement leads us to posit that the entire system $ABCE$ is in a pure state. Consequently, the von Neumann entropy of the combined system, $S(ABCE)$, is zero.
Before proceeding, we must clarify our notation. Throughout this work, the term $S(I|J)$ denotes the entropy of subsystem $I$ evaluated in the state, $\rho_{I|J}$, which results after a measurement on subsystem $J$. The entropy is calculated as $S(I|J):=S(\rho_{I|J})=-\text{Tr}[\rho_{I|J}\log_{2}\rho_{I|J}]$.
This definition should not be confused with the standard conditional von Neumann entropy, which is an algebraic property of the initial (pre-measurement) state $\rho_{IJ}$. The standard definition, which we denote as $S(I|J)_{\text{std}}$, is given by $S(I|J)_{\text{std}}=S(IJ)-S(J)$, or explicitly:
$S(I|J)_{\text{std}}=-\text{Tr}[\rho_{IJ}\log_{2}\rho_{IJ}]+\text{Tr}[\rho_{J}\log_{2}\rho_{J}]$ (1)
Crucially, our notation $S(I|J)$ refers to the entropy of a state produced by a physical measurement process, while the standard definition does not.
With this notational framework, we can establish the relationships needed to bound Eve’s information. Since the overall state of $ABCE$ is pure, the entropy of any subsystem is equal to that of its complement; thus, $S(ABC)=S(E)$. Furthermore, a measurement on subsystem $A$ collapses the remaining system $BCE$ into a new pure state. For any tripartite pure state, the entropy of a subsystem is equal to that of its complement. This physical principle directly leads to the equality $S(BC|A)=S(E|A)$, where both quantities are entropies of the respective subsystems in the post-measurement state conditioned on the outcome at $A$.
Eve’s knowledge of the measurement outcome of $A$ is upper-bounded by the Holevo quantity $\chi_{A}$, which is the initial entropy of Eve’s system reduced by the entropy remaining after Alice’s measurement. This can be written as $\chi_{A}=S(E)-S(E|A)$. Using the relations established above, this bound becomes:
$\chi_{A}=S(ABC)-S(BC|A).$ (2)
This framework can be extended to sequential key generation. Here, the analysis of a subsequent key between $B$ and $C$ is performed on the conditional state $\rho_{BC|A}$. This approach, where the protocol conditions on $A$’s data and the security proof assumes this data is known to all parties including Eve, is a deliberate choice made for two critical security reasons. First, the assumption that Eve has access to $A$’s data allows us to establish the most conservative (worst-case) upper bound on her information for the subsequent key. Second, by generating the key only from the data corresponding to the conditional state $\rho_{BC|A}$, we ensure the new key is statistically independent (uncorrelated) of $A$’s original data. This prevents vulnerabilities that could arise from cross-key correlations.
It is important to note that this conditioning does not require a sequential quantum measurement process. The protocol imposes no specific temporal order on the measurements; each party simply measures their respective quantum state upon receipt, and the ’conditioning’ is performed entirely during classical post-processing. As will be detailed in the key rates analysis (Sec. III.2), this classical step requires at least one of the parties ($B$ or $C$) to have knowledge of $A$’s data. Therefore, the analysis of this stage is not an examination of a hypothetical scenario, but a direct assessment of the cryptographic resource available to the remaining parties under this robust security model. For a key generated between $B$ and another party using this state, the information leaked to Eve is bounded by quantities that depend on the entropy of the full remaining system, $S(BC|A)$, reduced by the entropy of the system left after $B$’s measurement. The bounds are thus given by:
$\chi_{B|A}=S(BC|A)-S(C|AB)$ (3)
and similarly, for a key generated by $C$:
$\chi_{C|A}=S(BC|A)-S(B|AC).$ (4)
@@
$\Sigma_{ij}=\frac{1}{2}\left\langle\hat{x}_{i}\hat{x}_{j}+\hat{x}_{j}\hat{x}_{i}\right\rangle-\left\langle\hat{x}_{i}\right\rangle\left\langle\hat{x}_{j}\right\rangle,$ (5)
where $\left\langle\cdot\right\rangle$ denotes expectation values. Diagonal elements of the matrix indicate the variance of each quadrature, representing their uncertainties, while off-diagonal elements represent correlations between different quadratures.
@@
$S=\sum_{i}\left[\left(\nu_{i}+\frac{1}{2}\right)\log\left(\nu_{i}+\frac{1}{2}\right)-\left(\nu_{i}-\frac{1}{2}\right)\log\left(\nu_{i}-\frac{1}{2}\right)\right].$ (6)
@@
where $\Sigma_{p}$ denotes the covariance matrix of the unmeasured modes of the multipartite state, $\Sigma_{q}$ represents the covariance matrix of the measured modes, $\Sigma_{pq}$ is the covariance matrix between the measured and unmeasured modes, and $X$ is defined as $Diag(1,0)$ for measurements in the x-quadrature and $MP$ in the Moore-Penrose inverse.
@@
Throughout this work, we assume homodyne detection of the $x$-quadrature unless stated otherwise. For all graph states studied here, except the GHZ/W state, the quadratures are symmetric in phase space, meaning the variances of the $x$- and $p$-quadratures are equal. As a result, the choice of measurement quadrature does not affect the key rate or security analysis. In the case of the GHZ/W state, our construction yields lower variance in the $x$-quadrature, making it the favorable choice for measurement. We also considered heterodyne detection as an alternative, but it was consistently found to be suboptimal compared to homodyne detection across all scenarios analyzed. To maintain clarity and avoid redundancy, all references to measurement in this paper should be understood as $x$-quadrature homodyne detection.
III.2 Key Rates
Having addressed the complexities of multi-user QKD in the contexts of information reconciliation and the estimation of the Holevo bound, we are now positioned to establish expressions for various key rates [33, 34, 31]. For a three-user system, three distinct types of keys can be generated:
@@ III.2 Key Rates
$\beta$ is the reconciliation efficiency [35] and the superscript signifies the reference data of the user $i$ used to generate the key. The efficiency $\beta$ is a crucial practical parameter that accounts for the fact that real-world error correction protocols, used by the parties to reconcile their correlated data into an identical string, are not perfectly efficient and have a communication overhead.
@@
Once a CKA with user $i$ as the reference is complete, the conditional data possessed by the remaining two users, $j$ and $k$, can be utilized to generate a new, independent key. Following a successful CKA, both users $j$ and $k$ now possess the measurement outcomes of user $i$. This allows either party to act as the reference for the subsequent key generation. A secure key can be generated from this conditional resource provided that the conditional mutual information between $j$ and $k$, $I(j:k|i)$, is greater than the information Eve can obtain about the raw key.
To determine the Holevo bound for this subsequent stage, we apply the same security principles as before. We pessimistically assume that Eve also knows the measurement outcomes of $i$. The security analysis is then performed on the conditional state $\rho_{jk|i}$, where we posit that Eve holds a purification of this state. Following the arguments presented in Sec. III.1, we can then calculate the bound on Eve’s information for this specific conditional resource. The key rate for this scenario depends on which user acts as the reference for error correction. If user $j$ is the reference, the key rate is given by:
@@
A user can generate independent bipartite keys with the users with whom it is correlated, provided the mutual information of this user (say $i$) with the rest of the users is higher than the mutual information between the other users (ideally zero). The key rates are:
@@
and $K^{k}(i:k|j)$. In the scenario where the mutual information between the other users is zero, the equation simplifies to $K^{k}(i:k)$. Notice that even when $j$ and $k$ are identical, the order of key generation influences the key rates. This sequential dependence is critical. After user $i$ establishes a key with user $j$ (corresponding to the rate $K^{j}(i:j)$), user $i$ now possesses the measurement outcomes of user $j$ after successful error correction. This knowledge enables the subsequent step. For the key rate $K^{k}(i:k|j)$, where user $k$ provides the reference data, key generation is possible because user $i$ has the necessary conditional information (its own data and $j$’s data) to perform error correction and privacy amplification. However, alternate reconciliation with key rate $K^{i}(i:k|j)$ would not be described by this framework, as it would require user $k$ to perform error correction and privacy amplification, for which it would need access to $j$’s measurement outcomes—information they do not have. It is more useful to consider the sum of independent keys that $i$ can generate with other users. The actual key rates with each user depend on the information disclosed for error correction, specifically the syndrome length. The sum of independent keys is given by:
$\begin{split}K^{j,k}(i\!:\!jk)&=H(j,k)-\zeta H(j,k|i)-\chi_{jk}\\ &=K^{j}(i\!:\!j)+K^{k}(i\!:\!k|j)\end{split}$ (13)
As discussed in Sec. II, $H(j,k|i)$ represents the optimal amount of information that $j$ and $k$ need to communicate to $i$ to facilitate error correction and $\zeta\geq 1$ accounts for inefficient information reconciliation. For the symmetric case, the independent key rates can be computed as follows:
$K(i\!:\!j)=K(i\!:\!k)$ $=\tfrac{1}{2}\Big[H(j,k)-\zeta H(\mathbf{S}_{j},\mathbf{S}_{k})$
$\quad-S(ijk)+S(i\mid jk)\Big]$ (14)
@@
III.3 Finite Size Analysis